# Monotone Lipschitz-Gradient Denoiser: Explainability of Operator Regularization Approaches Free From Lipschitz Constant Control

### v2

@@ V-D Prior Art II: Other Related Works
Characterization of implicit regularizer: In [2], the softmax activation function is shown to be Moreau’s proximity operator of a certain convex function by using [39, Corollary 24.5], which is related to, but is different from, Theorem 1(b). Specifically, [39, Corollary 24.5] states that, given $\varphi\in\Gamma_{0}({\mathcal{H}})$ and $f\in\Gamma_{0}({\mathcal{H}})$ such that $\varphi=f-(1/2)\left\|\cdot\right\|^{2}$, it holds that ${\rm s\mbox{-}Prox}_{\varphi}=\nabla f^{*}$. Here, $\varphi$ is implicitly assumed to be $1$-strongly convex. Meanwhile, by changing $\psi^{*}$ in Theorem 1(b) by $f$, it can essentially be stated as follows: given $\varphi+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ and $f-(\beta/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ with $\varphi=f-(1/2)\left\|\cdot\right\|^{2}$, it holds that ${\rm s\mbox{-}Prox}_{\varphi}=\nabla f^{*}$. Here, $\varphi$ is $(1-\beta)$-weakly convex (while $f$ is $\beta$-strongly convex) in contrast to the former case.
@@
Proximity operator of weakly convex function in a linear inverse problem: In [52], the proximity operator (in the sense of Definition 4 essentially) of weakly convex functions has been studied in a linear inverse problem, and the convergence of the forward-backward splitting algorithm with the s-prox operator has been analyzed. While the main results therein are related to Theorem 2, the present study has a wider scope. Specifically, it includes Theorem 1, Propositions 2 and 3, and Theorem 3 as well as the discussions in Section IV-C to make the idea be applicable to other operator splitting algorithms. Moreover, the study in [52] considers the case in which (i) ${\mathcal{H}}$ is the Euclidean space, and (ii) the range of $\varphi$ is ${\mathbb{R}}$. whereas the present study concerns the case in which (i) ${\mathcal{H}}$ is a general Hilbert space, and (ii) $\varphi$ may take the value +$\infty$+ ∞.
@@ VI-A Firm Shrinkage Plugged into Algorithm 1
We consider the linear system model ${\boldsymbol{y}}:={\boldsymbol{A}}{\boldsymbol{x}}_{\diamond}+{\boldsymbol{\varepsilon}}\in{\mathbb{R}}^{m}$, where ${\boldsymbol{x}}_{\diamond}\in{\mathcal{H}}:={\mathbb{R}}^{n}$ is the $n$-dimensional Euclidean vector to estimate, ${\boldsymbol{A}}\in{\mathbb{R}}^{m\times n}$ is the sensing matrix, and ${\boldsymbol{\varepsilon}}\in{\mathbb{R}}^{m}$ is the noise vector. We consider the overdetermined case $\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})>0$, where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue. (See [51] for the case of $\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})=0$.) We suppose that ${\boldsymbol{x}}_{\diamond}$ is piecewise constant so that the difference vector ${\boldsymbol{D}}{\boldsymbol{x}}=[x_{1}-x_{2},x_{2}-x_{3},\cdots,x_{n-1}-x_{n}]^{\top}$ is sparse, where ${\boldsymbol{D}}:=[{\boldsymbol{I}}~{}{\boldsymbol{0}}]-[{\boldsymbol{0}}~{}{\boldsymbol{I}}]\in{\mathbb{R}}^{(n-1)\times n}$ with the identity matrix ${\boldsymbol{I}}$ and the zero vector ${\boldsymbol{0}}$ of length $n-1$. We therefore let $L:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n-1}=:\mathcal{U}:{\boldsymbol{x}}\mapsto{\boldsymbol{D}}{\boldsymbol{x}}$. The data fidelity term is set to $f({\boldsymbol{x}}):=(1/2)\left\|{\boldsymbol{A}}{\boldsymbol{x}}-{\boldsymbol{y}}\right\|^{2}_{2}$. In this case, since the Hessian matrices of $f$ and $\hat{f}$ are given by $\nabla^{2}f({\boldsymbol{x}})={\boldsymbol{A}}^{\top}{\boldsymbol{A}}$ and $\nabla^{2}\hat{f}({\boldsymbol{x}})={\boldsymbol{A}}^{\top}{\boldsymbol{A}}-(\rho/\left\|{\boldsymbol{D}}\right\|^{2}){\boldsymbol{D}}^{\top}{\boldsymbol{D}}$, respectively, at every ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, $f$ is $\rho$-strongly convex for $\rho:=\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})$, and $\hat{f}$ is $\kappa$-smooth for $\kappa:=\|{\boldsymbol{A}}^{\top}{\boldsymbol{A}}-(\rho/\left\|{\boldsymbol{D}}\right\|^{2}){\boldsymbol{D}}^{\top}{\boldsymbol{D}}\|$.
@@ VI-B Experimental Verification of Theorem 3
As $\tilde{f}:=\mu f-~{}^{\lambda_{2}}(\left\|\cdot\right\|_{1})\circ{\boldsymbol{D}}$ is a smooth convex function with its gradient available in a closed form, the minimization problem in (48) can be solved by the operator splitting algorithm such as the primal-dual splitting method with convergence guarantee.111111 This is possible in this specific example, because the function $\varphi$ corresponding to the firm shrinkage is available. In general, $\varphi$ could be unavailable, and thus this kind of approach cannot be used in such a case. To demonstrate the validity of Theorem 3 by simulation, we apply the Condat–Vũ algorithm (form II) to (48) to see whether Algorithm 1 converges to the same point. We show, more specifically, that the discrepancy is vanishing between $({\boldsymbol{x}}_{k},{\boldsymbol{u}}_{k})$ of Algorithm 1 and $(\tilde{{\boldsymbol{x}}}_{k},\tilde{{\boldsymbol{u}}}_{k})$ of the Condat–Vũ algorithm applied to (48), where “discrepancy” is quantified as follows:
@@
We let $n:=256$, $m=1024$, and ${\boldsymbol{x}}_{\diamond}$ depicted in Fig. 3 is used. For Algorithm 1, we set $\delta:=1$, $\tau:=0.9(\sigma\left\|{\boldsymbol{D}}\right\|^{2}+\kappa/2)^{-1}$, $\lambda_{1}:=2.5$, and $\lambda_{2}:=5$. We repeat that $\lambda_{1}$ does not change the solution, see (47). For the approach based on (48), we set $\sigma:=0.2$, and all other parameters are the same as for Algorithm 1. Figure 4 plots the discrepancy given in (49) (which is labeled as ‘${\boldsymbol{x}},{\boldsymbol{u}}$’); the discrepancy in terms of the primal variables and that in terms of the dual variables are also plotted, labeled as ‘${\boldsymbol{x}}$’ and ‘${\boldsymbol{u}}$’, respectively. It can be seen that the discrepancy vanishes as time goes by.
For a comparison, we also test a straightforward approach (which has no theoretical guarantee): apply the Condat-Vũ algorithm (form II) directly to the nonconvex formulation (46), which is essentially the same as (48). Here, ${\rm s\mbox{-}Prox}_{\sigma g^{*}}$ appearing in the algorithm is replaced121212 This replacement is heuristic, and there is no theoretical support for that. by ${\rm Id}-\sigma{\rm s\mbox{-}Prox}_{\varphi_{\lambda_{2}}^{\rm MC}/(\mu\sigma)}\circ(\sigma^{-1}{\rm Id})$, where ${\rm s\mbox{-}Prox}_{\varphi_{\lambda_{2}}^{\rm MC}/(\mu\sigma)}$ is the proximity operator in the sense of Definition 4, which coincides with the firm shrinkage with $\lambda_{1}:=1/(\mu\sigma)$. The $\sigma$ here needs to be chosen so that $\lambda_{1}<\lambda_{2}$. The results are plotted in Fig. 5. It can be seen that the discrepancy (‘${\boldsymbol{x}},{\boldsymbol{u}}$’) does not vanish. Although the discrepancy in terms of ‘${\boldsymbol{x}}$’ vanishes, there is no theoretical guarantee for that, to the best of authors’ knowledge.
@@ Appendix A Proof of Proposition 1
$p\!\in\!\mathbf{Prox}_{f}(x)\Leftrightarrow 0\!\in\!\partial J_{x}(p)\Leftrightarrow x\!\in\!\partial J_{0}(p)\Leftrightarrow p\!\in\!\big{(}\partial J_{0}\big{)}^{-1}(x),$
@@
For contradiction, suppose, for an arbitrarily fixed $x\in{\mathcal{H}}$, that $\mathbf{Prox}_{f}(x)$ contains two distinct vectors $T(x)$ and $T(x)+\delta$ for some $\delta\in{\mathcal{H}}\setminus\{0\}$. The monotonicity of $\mathbf{Prox}_{f}$ suggests that, for every ϵ++italic-$\epsilon\in{\mathbb{R}}_{++}$, it holds that
@@ Appendix B Preservation of averaged nonexpansiveness
Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$. Then, for every $\alpha\in(0,1)$ and ϱ++italic-$\varrho\in{\mathbb{R}}_{++}$, the following hold.
@@
    $\varrho T$ is $\alpha$-averaged nonexpansive if and only if $T\circ(\varrho{\rm Id})$ is $\alpha$-averaged nonexpansive. In particular, $T$ is $\varrho$-cocoercive if and only if $T\circ(\varrho{\rm Id})$ is firmly ((1/2)-averaged) nonexpansive.
@@
$\hskip 80.00012pt\exists N:{\mathcal{H}}\rightarrow{\mathcal{H}}\mbox{ nonexpansive.}$. (B.3)
@@ Appendix C Proof of Example 1
[…]
@@
(${}^{\lambda}(h\circ\left\|\cdot\right\|)(x)=\min_{y\in{\mathcal{H}}}\Big{(}h(\left\|y\right\|)+\frac{1}{2\lambda}\left\|y-x\right\|^{2}\Big{)}.$ (C.4)
If $x=0$, (${}^{\lambda}(h\circ\left\|\cdot\right\|)(0)=~{}^{\lambda}(h\circ\left|\cdot\right|)(\left\|0\right\|)=h(0)$. We now assume that $x\neq 0$. Based on the orthogonal decomposition of ${\mathcal{H}}=\mathcal{M}\oplus\mathcal{M}^{\perp}$ with $\mathcal{M}:={\rm span}\{x\}$ and its orthogonal complement $\mathcal{M}^{\perp}$, every $y$ can be represented by $y=\xi x+y_{\perp}$ with $\xi\in{\mathbb{R}}$ and $y_{\perp}\in\mathcal{M}^{\perp}$. Thus, by the Pythagorean theorem, the Moreau envelope in (C.4) can be represented as follows:
($~{}^{\lambda}(h\circ\left\|\cdot\right\|)(x)=$ $~{}\min_{\xi\in{\mathbb{R}},y_{\perp}\in\mathcal{M}^{\perp}}h\Big{(}\sqrt{\left\|\xi x\right\|^{2}+\left\|y_{\perp}\right\|^{2}}\Big{)}$
@@
$=$ ($~{}^{\lambda}(h\circ\left|\cdot\right|)(\left\|x\right\|),$ (C.5)
@@
Proof of Example 1: Letting $h:x\mapsto x$ and $\lambda:=\lambda_{2}$ in Lemma C.1 gives ${}^{\lambda_{2}}\left\|\cdot\right\|(x)=~{}^{\lambda_{2}}\left|\cdot\right|(\left\|x\right\|)$. This, together with $(\left|\cdot\right|)_{\lambda_{2}}=\varphi_{\lambda_{2}}^{\rm MC}$, verifies that $(\left\|\cdot\right\|)_{\lambda_{2}}=\varphi_{\lambda_{2}}^{\rm MC}\circ\left\|\cdot\right\|$. Recall now that the MC function given in (38) is a symmetric nondecreasing function. Hence, letting $h=\varphi_{\lambda_{2}}^{\rm MC}$ and $\lambda:=\lambda_{1}$ in Lemma C.1 yields ${\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}\circ\left\|\cdot\right\|}(x)=(x/\left\|x\right\|){\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}\circ\left|\cdot\right|}(\left\|x\right\|)=(x/\left\|x\right\|){\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}}(\left\|x\right\|)=(x/\left\|x\right\|){\rm firm}_{\lambda_{1},\lambda_{2}}(\left\|x\right\|)$, where the second equality is owing to the symmetry of $\varphi_{\lambda_{2}}^{\rm MC}$. See Section V-A and also [17, 74, 29] for the third equality. $\blacksquare$

### v3

@@ V-D Prior Art II: Other Related Works
Characterization of implicit regularizer: In [2], the softmax activation function is shown to be Moreau’s proximity operator of a certain convex function by using [39, Corollary 24.5], which is related to, but is different from, Theorem 1(b). Specifically, [39, Corollary 24.5] states that, given $\varphi\in\Gamma_{0}({\mathcal{H}})$ and $f\in\Gamma_{0}({\mathcal{H}})$ such that $\varphi=f-(1/2)\left\|\cdot\right\|^{2}$, it holds that ${\rm s\mbox{-}Prox}_{\varphi}=\nabla f^{*}$. Here, $f$ is implicitly assumed to be $1$-strongly convex. Meanwhile, by changing $\psi^{*}$ in Theorem 1(b) by $f$, it can essentially be stated as follows: given $\varphi+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ and $f-(\beta/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$ with $\varphi=f-(1/2)\left\|\cdot\right\|^{2}$, it holds that ${\rm s\mbox{-}Prox}_{\varphi}=\nabla f^{*}$. Here, $\varphi$ is $(1-\beta)$-weakly convex (while $f$ is $\beta$-strongly convex) in contrast to the former case.
@@
Proximity operator of weakly convex function in a linear inverse problem: In [52], the proximity operator (in the sense of Definition 4 essentially) of weakly convex functions has been studied in a linear inverse problem, and the convergence of the forward-backward splitting algorithm with the s-prox operator has been analyzed. While the main results therein are related to Theorem 2, the present study has a wider scope. Specifically, it includes Theorem 1, Propositions 2 and 3, and Theorem 3 as well as the discussions in Section IV-C to make the idea be applicable to other operator splitting algorithms. Moreover, the study in [52] considers the case in which (i) $\mathcal{H}$} is the Euclidean space, and (ii) the range of $\varphi$ is $\mathbb{R}$}. whereas the present study concerns the case in which (i) $\mathcal{H}$} is a general Hilbert space, and (ii) $\varphi$ is allowed to take the value $\infty$.
@@ VI-A Firm Shrinkage Plugged into Algorithm 1
We consider the linear system model ${\boldsymbol{y}}:={\boldsymbol{A}}{\boldsymbol{x}}_{\diamond}+{\boldsymbol{\varepsilon}}\in{\mathbb{R}}^{m}$, where ${\boldsymbol{x}}_{\diamond}\in{\mathcal{H}}:={\mathbb{R}}^{n}$ is the $n$-dimensional Euclidean vector to estimate, ${\boldsymbol{A}}\in{\mathbb{R}}^{m\times n}$ is the sensing matrix, and ${\boldsymbol{\varepsilon}}\in{\mathbb{R}}^{m}$ is the noise vector. We consider the overdetermined case $\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})>0$, where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue. (See [51] for the case of $\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})=0$.) We suppose that $\boldsymbol{x}$$_{\diamond}$ is piecewise constant so that the difference vector ${\boldsymbol{D}}{\boldsymbol{x}}=[x_{1}-x_{2},x_{2}-x_{3},\cdots,x_{n-1}-x_{n}]^{\top}$ is sparse, where ${\boldsymbol{D}}:=[{\boldsymbol{I}}~{\boldsymbol{0}}]-[{\boldsymbol{0}}~{\boldsymbol{I}}]\in{\mathbb{R}}^{(n-1)\times n}$ with the identity matrix $\boldsymbol{I}$} and the zero vector $\boldsymbol{0}$} of length n1n-1. We therefore let $L:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n-1}=:\mathcal{U}:{\boldsymbol{x}}\mapsto{\boldsymbol{D}}{\boldsymbol{x}}$. The data fidelity term is set to $f({\boldsymbol{x}}):=(1/2)\left\|{\boldsymbol{A}}{\boldsymbol{x}}-{\boldsymbol{y}}\right\|^{2}_{2}$. In this case, since the Hessian matrices of $f$ and $\hat{f}$ are given by $\nabla^{2}f({\boldsymbol{x}})={\boldsymbol{A}}^{\top}{\boldsymbol{A}}$ and $\nabla^{2}\hat{f}({\boldsymbol{x}})={\boldsymbol{A}}^{\top}{\boldsymbol{A}}-(\rho/\left\|{\boldsymbol{D}}\right\|^{2}){\boldsymbol{D}}^{\top}{\boldsymbol{D}}$, respectively, at every ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, $f$ is $\rho$-strongly convex for $\rho:=\lambda_{\min}({\boldsymbol{A}}^{\top}{\boldsymbol{A}})$, and $\hat{f}$ is $\kappa$-smooth for $\kappa:=\|{\boldsymbol{A}}^{\top}{\boldsymbol{A}}-(\rho/\left\|{\boldsymbol{D}}\right\|^{2}){\boldsymbol{D}}^{\top}{\boldsymbol{D}}\|$.
@@ VI-B Experimental Verification of Theorem 3
As $\tilde{f}:=\mu f-~^{\lambda_{2}}(\left\|\cdot\right\|_{1})\circ{\boldsymbol{D}}$ is a smooth convex function with its gradient available in a closed form, the minimization problem in (48) can be solved by the operator splitting algorithm such as the primal-dual splitting method with convergence guarantee. [^11] This is possible in this specific example, because the function $\varphi$ corresponding to the firm shrinkage is available. In general, $\varphi$ could be unavailable, and thus this kind of approach cannot be used in such a case. To demonstrate the validity of Theorem 3 by simulation, we apply the Condat–Vũ algorithm (form II) to (48) to see whether Algorithm 1 converges to the same point. We show, more specifically, that the discrepancy is vanishing between $({\boldsymbol{x}}_{k},{\boldsymbol{u}}_{k})$ of Algorithm 1 and $(\tilde{{\boldsymbol{x}}}_{k},\tilde{{\boldsymbol{u}}}_{k})$ of the Condat–Vũ algorithm applied to (48), where “discrepancy” is quantified as follows:
@@
We let $n:=256$, $m=1024$, and $\boldsymbol{x}$$_{\diamond}$ depicted in Fig. 3 is used. For Algorithm 1, we set $\delta:=1$, $\tau:=0.9(\sigma\left\|{\boldsymbol{D}}\right\|^{2}+\kappa/2)^{-1}$, $\lambda_{1}:=2.5$, and $\lambda_{2}:=5$. We repeat that $\lambda_{1}$ does not change the solution, see (47). For the approach based on (48), we set $\sigma:=0.2$, and all other parameters are the same as for Algorithm 1. Figure 4 plots the discrepancy given in (49) (which is labeled as ${\boldsymbol{x}},{\boldsymbol{u}}$’); the discrepancy in terms of the primal variables and that in terms of the dual variables are also plotted, labeled as $\boldsymbol{x}$}’ and $\boldsymbol{u}$}’, respectively. It can be seen that the discrepancy vanishes as time goes by.
For a comparison, we also test a straightforward approach (which has no theoretical guarantee): apply the Condat-Vũ algorithm (form II) directly to the nonconvex formulation (46), which is essentially the same as (48). Here, ${\rm s\mbox{-}Prox}_{\sigma g^{*}}$ appearing in the algorithm is replaced [^12] This replacement is heuristic, and there is no theoretical support for that. by ${\rm Id}-\sigma{\rm s\mbox{-}Prox}_{\varphi_{\lambda_{2}}^{\rm MC}/(\mu\sigma)}\circ(\sigma^{-1}{\rm Id})$, where ${\rm s\mbox{-}Prox}_{\varphi_{\lambda_{2}}^{\rm MC}/(\mu\sigma)}$ is the proximity operator in the sense of Definition 4, which coincides with the firm shrinkage with $\lambda_{1}:=1/(\mu\sigma)$. The $\sigma$ here needs to be chosen in such a way that $\lambda_{1}<\lambda_{2}$. The results are plotted in Fig. 5. It can be seen that the discrepancy (${\boldsymbol{x}},{\boldsymbol{u}}$’) does not vanish. Although the discrepancy in terms of $\boldsymbol{x}$}’ vanishes, there is no theoretical guarantee for that, to the best of authors’ knowledge.
@@ Appendix A Proof of Proposition 1
$\hskip-20.00003ptp\!\in\!\mathbf{Prox}_{f}(x)\Leftrightarrow 0\!\in\!\partial J_{x}(p)\Leftrightarrow x\!\in\!\partial J_{0}(p)\Leftrightarrow p\!\in\!\big(\partial J_{0}\big)^{-1}(x),$
@@
For contradiction, suppose, for an arbitrarily fixed $x\in{\mathcal{H}}$, that $\mathbf{Prox}_{f}(x)$ contains two distinct vectors $T(x)$ and $T(x)+\delta$ for some $\delta\in{\mathcal{H}}\setminus\{0\}$. The monotonicity of $\mathbf{Prox}_{f}$ suggests that, for every $\epsilon\in{\mathbb{R}}_{++}$, it holds that
@@ Appendix B Preservation of averaged nonexpansiveness
Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$. Then, for every $\alpha\in(0,1)$ and $\varrho\in{\mathbb{R}}_{++}$, the following hold.
@@
    $\varrho T$ is $\alpha$-averaged nonexpansive if and only if $T\circ(\varrho{\rm Id})$ is $\alpha$-averaged nonexpansive. In particular, $T$ is $\varrho$-cocoercive if and only if $T\circ(\varrho{\rm Id})$ is firmly (1/2-averaged) nonexpansive.
@@
$\hskip 80.00012pt\exists N:{\mathcal{H}}\rightarrow{\mathcal{H}}\mbox{ nonexpansive.}$ (B.3)
@@ Appendix C Proof of Example 1
Let $\left\|\cdot\right\|$ be a norm in a real Hilbert space $\mathcal{H}$}, and $h:{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ be a nonincreasing function; i.e., ${\mathbb{R}}_{+}\ni a\leq b\in{\mathbb{R}}_{+}\Rightarrow h(a)\leq h(b)$. Assume that the s-prox operator of $h$ is well-defined. Then, for every $\lambda\in{\mathbb{R}}_{++}$ and every $x\in{\mathcal{H}}$, it holds that $^{\lambda}(h\circ\left\|\cdot\right\|)(x)=~^{\lambda}(h\circ\left|\cdot\right|)(\left\|x\right\|)$ and ${\rm s\mbox{-}Prox}_{\lambda h\circ\left\|\cdot\right\|}(x)=\begin{cases}\dfrac{{\rm s\mbox{-}Prox}_{\lambda h\circ\left|\cdot\right|}(\left\|x\right\|)}{\left\|x\right\|}x&\mbox{if }x\neq 0\\ 0&\mbox{if }x=0.\end{cases}$
@@
$^{\lambda}(h\circ\left\|\cdot\right\|)(x)=\min_{y\in{\mathcal{H}}}\Big(h(\left\|y\right\|)+\frac{1}{2\lambda}\left\|y-x\right\|^{2}\Big).$ (C.4)
If $x=0$, $^{\lambda}(h\circ\left\|\cdot\right\|)(0)=~^{\lambda}(h\circ\left|\cdot\right|)(\left\|0\right\|)=h(0)$. We now assume that $x\neq 0$. Based on the orthogonal decomposition of $\mathcal{H}$$\mathcal{M}\oplus\mathcal{M}^{\perp}$ with $\mathcal{M}:={\rm span}\{x\}$ and its orthogonal complement $\mathcal{M}^{\perp}$, every $y$ can be represented by $y=\xi x+y_{\perp}$ with $\xi\in{\mathbb{R}}$ and $y_{\perp}\in\mathcal{M}^{\perp}$. Thus, by the Pythagorean theorem, the Moreau envelope in (C.4) can be represented as follows:
$^{\lambda}(h\circ\left\|\cdot\right\|)(x)=$ $~\min_{\xi\in{\mathbb{R}},y_{\perp}\in\mathcal{M}^{\perp}}h\Big(\sqrt{\left\|\xi x\right\|^{2}+\left\|y_{\perp}\right\|^{2}}\Big)$
@@
$=$ $^{\lambda}(h\circ\left|\cdot\right|)(\left\|x\right\|),$ (C.5)
@@
Proof of Example 1: We only prove the case of $x\neq 0$, as the case of $x=0$ is straightforward. Letting $h:x\mapsto x$ and $\lambda:=\lambda_{2}$ in Lemma C.1 gives $^{\lambda_{2}}\left\|\cdot\right\|(x)=~^{\lambda_{2}}\left|\cdot\right|(\left\|x\right\|)$. This, together with $(\left|\cdot\right|)_{\lambda_{2}}=\varphi_{\lambda_{2}}^{\rm MC}$, verifies that $(\left\|\cdot\right\|)_{\lambda_{2}}=\varphi_{\lambda_{2}}^{\rm MC}\circ\left\|\cdot\right\|$. Recall now that the MC function given in (38) is a symmetric nondecreasing function. Hence, letting $h=\varphi_{\lambda_{2}}^{\rm MC}$ and $\lambda:=\lambda_{1}$ in Lemma C.1 yields ${\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}\circ\left\|\cdot\right\|}(x)=(x/\left\|x\right\|){\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}\circ\left|\cdot\right|}(\left\|x\right\|)=(x/\left\|x\right\|){\rm s\mbox{-}Prox}_{\lambda_{1}\varphi_{\lambda_{2}}^{\rm MC}}(\left\|x\right\|)=(x/\left\|x\right\|){\rm firm}_{\lambda_{1},\lambda_{2}}(\left\|x\right\|)$, where the second equality is owing to the symmetry of $\varphi_{\lambda_{2}}^{\rm MC}$. See Section V-A and also [17, 74, 29] for the third equality. $\blacksquare$